The area of a rectagle is 357cm2. The width of the rectangle is 4cm less than the length. What are the dimensions of the rectangle
Answer by checkley77(12844)

W=L-4
AREA=WL
357=(L-4)L
357=L^2-4L
L^2-4L-357=0
(L-21)(L+17)=357
L-21=0
L=21 ANS. FOR THE LENGTH.
W=21-4=17 ANS. FOR THE WIDTH.
PROOF:
357=21*17
357=357